A line passes through the points -2 and -7 and has a slope of 3 write an equation for this line
Answer by mananth(16946)


the general form of equation is
(y-y1) = m(x-x1)
m = slope
and (x1,y1) are the co ordinates of a point on the line
(y-(-7)) = 3(x-(-2))
y+7= 3x+6
3x-y=1